OpenStudy (anonymous):

Can someone help with this question? A candle in the shape of a circular cone has a base of radius r and height of h that is the same length as the radius. Which expresses the ratio of the volume of the candle to its surface area (including the base) for cone v=1/3 pi r^2 and sa= pi r^2 + pi r sqrt r^2+h^2

OpenStudy (anonymous):

Volume of a cone =1/3 pi r^3 Surface Area of cone = (pi *r*l +pi r^2)

OpenStudy (anonymous):